# Experimental identification of topological defects in 2D colloidal glass

### v1

@@ Abstract
Topological defects are singularities in the order parameter space that are mathematically described by topological invariants and cannot be removed by continuous transformations. These defects play a significant role in various fields, ranging from cosmology to solid-state physics and biological matter. The definition of these irregularities requires an ordered reference configuration, leading to decades of debate about the existence of topological defects in disordered systems, such as glasses. Recently, it has been proposed that well-defined topological defects might emerge in the dynamical properties of glasses under deformation, potentially relating to their plastic behavior. In this study, we investigate a two-dimensional colloidal glass system composed of particles interacting by an effective magnetic repulsion. We reveal the presence of topological defects in the eigenspace of the vibrational frequencies of this experimental 2D amorphous solid. The vibrational density of states of this 2D amorphous material exhibits distinct glassy properties such as the presence of a boson peak anomaly. Through extensive numerical and theoretical quantitative analysis, we establish a robust positive correlation between the vibrational characteristics and the total number of topological defects. We furthermore show that defects of opposite charge tend to pair together and prove their local statistical correlation with the “soft spots”, the regions more prone to plastic flow. This work provides the experimental confirmation for the existence of topological defects in disordered systems revealing a complex interplay between topology, disorder, and vibrational behavior.
@@ Introduction
Topological defects (TD) represent a ubiquitous hallmark of nature across different scales and they are generally defined as singularities in a local order parameter [1]. They can appear in a wide range of physical systems [2, 3, 4], including but not limited to liquid crystals, superconductors, superfluids, ferromagnets, biological systems [5], and also our early universe [6]. Despite being microscopic in nature, TD can exert macroscopic influences on the behavior of the entire system. For example, they have a strong influence on the optical properties of liquid crystals and play crucial roles in various biological processes, including cell division, tissue formation, and the motion of cellular aggregates [7].
@@
In the realm of condensed matter physics [8], the emergence of these defects disrupts ordered states [9] influencing collective excitations and even triggering phase transitions, such as the Kosterlitz, Thouless, Halperin, Nelson, Young (KTHNY) transition [10, 11, 12] in two-dimensional solids, investigated in a colloidal mono-layer very similar to the given one [13]. These defects possess distinct (topological) charges and exhibit robustness against continuous structural deformations, rendering them essential for understanding the fundamental properties of materials and being totally different from other non-topological defects such as vacancies or interstitials.
@@
Topological defects in 3D crystalline solids manifest as dislocations and disclination lines [14, 9], providing structural irregularities within the otherwise ordered lattice, and the elementary carriers of plasticity. Dislocations, in particular, allow for the motion of crystalline planes sliding over each other (glide motion). The plasticity of crystalline solids is by now fully understood in terms of dislocation dynamics and dislocation networks, based on seminal concepts introduced by Taylor [15], Polanyi [16] and Orowan [17]. In 2D, topological defects are point defects and their thermal dissociation cause the elastic moduli to disappear [18].
@@
In amorphous solids, instead, the absence of long-range order complicates the identification and characterization of topological defects (since, by definition, an ordered background is needed in order to detect its irregularities), posing a big challenge in linking structure with dynamics and predicting mechanical properties from the undeformed material configuration. The chase for structural topological defects in amorphous solids, borrowing from the concepts widely used for crystals, is certainly not new and dates back to the early 70’s [19]. Despite various efforts [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], it was nevertheless concluded that topological arguments cannot be properly applied to the case of glasses and that the predictions of the dynamical and topological arguments disagree with results found in computer simulations [31].
@@
Therefore, apart from isolated cases [32, 33], the identification of the plastic carriers and of the “soft spots”, the regions (or particle clusters) more prone to plastic flow, in glasses remained based until now on the so-called structural indicators [34] (e.g., [35, 36, 37, 38, 39, 40, 41]) and on the phenomenological concept of shear transformation zones (STZ) [42, 43]. On the other hand, structural “soft spots”, where particle rearrangements are initiated, have been postulated to be analogous to dislocations in crystalline solids [44] but lacking any formal definition and topological character.
@@
Recent advances in simulation and theory [45] have pursued the idea of looking for topological defects in glasses using dynamical quantities such as the displacement vector field or the eigenvector field of normal mode vibrations, instead of the seemingly featureless structure. Following this idea, Baggioli et al. [46] have unveiled well-defined topological defects in polymer glasses under shear deformation. These topological defects are predicted by theory [47] as singularities in the microscopic displacement field, and are mathematically described by a continuous-valued Burgers vector (in accord with an early intuition of Kleman and Friedel [48]). In amorphous solids, the atomic displacement field under an external deformation comprises two vector fields, one of which is fully ordered and is called the affine displacement field. This is simply the trajectory along which each atom moves as a consequence of the external strain field. The other vector field is known as the nonaffine displacement field and is much more random and irregular. It arises due to the local force imbalance on each atom in its affine position, where it is subject to a non-vanishing force (due to the lack of inversion symmetry) communicated by its neighbours [49, 50]. In this sense, the affine component of the displacement field represents the ordered background disrupted by the topological defects living within (but not being equivalent to) the nonaffine displacements.
@@
Even more recently, mixing the ideas of [46] and [51], numerical investigations by Falk and collaborators [52] have identified saddle-like topological defects with quantized -$1$ charge in the displacement field, generating Eshelby-like quadrupolar fields which align to form the shear bands responsible for the yielding of amorphous solids [53, 54]. Interestingly, shear banding is directly related to the percolation of STZ, whose dynamics have been shown [55] to be intimately connected to vortex-like structures as those proposed by Wu et al.[51]. Finally, although topological charge neutrality is always enforced by the conservation of global elastic dipole charge, the yielding process in amorphous systems has been related to the clustering of net negative topological charge [56].
Topological defects in crystals can be also defined in terms of geometrical properties (curvature and torsion) [9]. A similar approach based on geometry has been recently proposed by Moshe, Procaccia and collaborators as the origin of plastic screening in amorphous materials [57, 58], where the defects are identified with the elastic (dipolar and quadrupolar) charges. Interesting results in this context can be also obtained in terms of a tensorial form of electromagnetism known as vector charge tensor gauge theory [59] that is able to reproduce the structure of static stress correlations in granular materials [60].
@@
Colloidal glasses have been proved to be an excellent experimental setup to test the validity of various measures and theories related to structure and dynamics [61, 62]. Using optical microscopy, one has direct access to the structural information at the particle level. In this experimental study, we use a specific type of a colloidal particles which become magnetised in the presence of a magnetic field and form a two-dimensional disordered structure. The particles interact via dipole-dipole repulsive interaction, so, based on particle positions, we can also directly infer the pairwise interaction energy. Using this information, we construct the dynamical matrix (Hessian) of the system, and perform the identification of topological defects in the field of normal modes. In this process, the analysis of the experimental data reveal unique features of vibrational characteristics of the system, and also the corresponding correlation with topology and “soft spots”. Our experimental study demonstrates the existence of well-defined topological defects in the vibrational field of a finite temperature 2D colloidal glass, suggesting further experimental studies that can employ these defects to understand the thermal and mechanical properties of many complex, disordered systems.
@@
Figure 1: a A representative snapshot of the colloidal system showcasing large and small particles colored in silver and pink, respectively. b The arrangement of colloidal particles forming a two-dimensional monolayer, subject to a constant external magnetic field $H$ oriented perpendicular to the surface.
Our experimental setup comprises a colloidal monolayer, where individual particles sediment to a flat water/air interface in a hanging-droplet configuration under the influence of gravity. These colloidal particles have two different masses and diameters, forming a binary mixture (Fig.1b), and exhibit Brownian motion within the two-dimensional plane (see Methods for details). The particle interactions are paramagnetic in nature, hence a magnetic dipole moment is induced in each particle when an external magnetic field is applied perpendicular to the surface, leading to a pairwise dipole-dipole interaction potential of the form $E_{\text{pot}}\sim 1/r^{3}$, where $r$ is the inter-particle distance. The strength of the interaction can be controlled by the external magnetic field $H$, setting the characteristic energy scale of the system. A crucial parameter governing the system’s behavior is represented by $\Gamma=E_{\text{pot}}/E_{\text{kin}}$, denoting the ratio of potential energy (stemming from this mutual dipolar interaction among particles) to the kinetic energy ($E_{\text{kin}}\sim T$) arising from thermal motion. It can be interpreted as an inverse temperature ore more didactically as dimensionless pressure. For our study, we set $\Gamma=423$, a value at which the system is deep in a glassy state [63]. Optical microscopy is employed to record the particle positions in the field of view at a certain interval of time (Fig.1a). The experimental setup is described in detail in [64], while the structure and dynamics of this system have been studied in previous work [65, 66].
@@
Using the long-range dipole-dipole interaction potential among the particles, we numerically construct the dynamical (Hessian) matrix which is diagonalized to obtain the eigenvectors $\vec{e}_{l}$ and eigenfrequencies $\lambda_{l}$. Here, $\lambda_{l}$ indicate the l-th eigenvalue with $l=1,...,2N$; $N$ being the number of particles in the field of view. From the eigenvalues, we derive the corresponding eigenfrequencies $\omega_{l}^{2}=\lambda_{l}$ (using a mass-rescaled Hessian, see Methods). As expected from amorphous solids at finite temperature [67], the spectrum displays a fraction of negative eigenvalues, corresponding to unstable modes with purely imaginary frequency. These unstable modes correspond to dynamics over regions of the potential landscape with locally negative curvature (e.g., saddle points and potential barriers) and they are widely observed in supercooled liquids as well [68]. For the unstable part of the spectrum ($\lambda<0$), we follow the practice of re-defining a positive definite frequency $\tilde{\omega}=-i\sqrt{\lambda}$ and plotting the corresponding vDOS on the negative frequency axes upon identifying $\tilde{\omega}=-\omega$. The results for the experimental vibrational density of states (vDOS) $D(\omega)$ are shown in Fig. 2a.
@@
As highlighted in the inset of Fig. 2a , the low frequency behavior of the vDOS shows a linear scaling, $D(\omega)\sim\omega$. This behavior is compatible with Debye’s law in two spatial dimensions but it is also a characteristic feature of the vDOS of systems with unstable modes [69, 70]. Interestingly, we observe that the vDOS is symmetric at low frequency and the linear scaling (including the corresponding slope) is the same for the stable and unstable branches. We also notice that only the stable branch can be directly interpreted as pertaining to vibrational modes stricto sensu.
@@
In Fig. 2b, we present the same vDOS normalized by the low-frequency linear (Debye) scaling, $D(\omega)/\omega$. In this representation, the presence of an anomalous peak is clear around a characteristic frequency $\omega_{\rm BP}$, marked within error bars by the shaded vertical strip ($15-17$ Hz). This vibrational anomaly over Debye’s law is known as “boson peak” and it is a common feature of amorphous materials [71]. Despite the huge effort devoted to characterize and comprehend this feature in 3D systems, its experimental detection in quasi two-dimensional (2D) amorphous materials remains limited [72, 73, 74]. In this respect, our result provides another experimental identification of the BP anomaly in a specific type of 2D amorphous materials – colloidal glasses – using directly the Hessian matrix rather than the covariance matrix.
@@
To measure the extent of mode localization, we also calculate the participation ratio $P(\omega)$, plotted in Fig. 2c. We observe a sharp peak corresponding to acoustic phonon-type excitations at very low (positive) frequency, and a secondary broader peak at higher frequency possibly linked to optical-like excitations due to the particle size mismatch. This secondary peak then drops sharply in correspondence of the Anderson-localized highest frequency modes which terminate at the Debye frequency of the system. The participation ratio for the unstable branch features the typical behavior observed many times in the past for glassy materials [75, 76].
@@
Figure 2: a The vibrational density of states $D(\omega)$ is presented, with the inset highlighting the low-frequency behavior. The dashed line represents a linear trend, consistent with Debye’s prediction $D(\omega)\sim\omega^{d-1}$, where $d=2$. b $D(\omega)/\omega$ versus $\omega$ is plotted at low frequencies, with shaded regions indicating the occurrence of the boson peak. c The participation ratio $P(\omega)$ versus $\omega$ is depicted, providing insights into the localization properties of vibrational modes across different frequency ranges.
@@
Figures 3a-d provide visual representations of the eigenvector fields at various frequencies, with filled red and blue circles denoting the locations of vortices and anti-vortices, respectively. Additionally, the color map indicates the local angle of the eigenvector field with respect to the $x$-axes. The structure of the eigenvector field is evidently highly heterogeneous, displaying several vortex-like structures and singularities. Notably, the eigenvectors at lower frequencies exhibit a smaller number of topological defects. This suggests that the number of defects grows with frequency and becomes more uniformly distributed in space. Significantly, at lower frequencies, there are fewer topological defects, and a rich, co-operative and swirling eigenfield structure. This might be related to the fact that at low-frequency the dynamics are dominated by plane waves, displaying a periodic structure of swirling (see e.g. Fig.3 in [77]). Conversely, at higher frequencies, there is a notable increase in defect density, with defects uniformly distributed throughout the space and reduced coherence in the eigenvector field. Interestingly, topological defects can be also found in the eigenvectors corresponding to unstable modes, e.g. panel a in Fig. 3.
Figure 4a presents the vibrational density of states $D(\omega)$ and the total number of defects $N_{d}$ against vibrational frequencies within the same frame. This highlights a robust correlation between these quantities across the entire frequency spectrum. The inset of Fig. 4a delves into the low-frequency behavior of $N_{d}$, where dashed line signifies the linear dependence on $\omega$. This linear dependence is different from what observed in a zero temperature simulation glass model in [51], where $N_{d}$ scales quadratically with the frequency at low energy. In the Supplementary Information, using simulations, we prove that this difference is due to the finite temperature dissipative effects and we recover the results of [51] in the zero temperature limit.
To further establish the correlation between vibrational properties and topological defects, we calculate the Pearson correlation coefficient, $C(\omega)$, at various $\omega$ values, considering data points within a specified frequency range $\omega-\Delta\omega/2$ to $\omega+\Delta\omega/2$. Figure 4b illustrates the variation of $C(\omega)$ for different frequency width values ($\Delta\omega=50$, $75$, and $100$). We observe that the fluctuations of $C(\omega)$ decrease as the frequency width increases. The proximity of $C$ values to unity across the majority of the spectrum signifies a robust correlation between $D(\omega)$ and $N_{d}$. However, intriguingly, instances of strong anti-correlation are observed within specific frequency ranges, as highlighted by gray vertical stripes in Fig. 4a and Fig. 4b. Particularly noteworthy is the presence of such anti-correlation within the frequency range $\omega\in[250,340]$, prompting further theoretical investigation to unravel its origin and its consequential impact on the system’s structure and dynamics.
@@
Figure 3: a-d Eigenvector fields within a rectangular field of view ($54a\times 40a$, where $a=20.78\leavevmode\nobreak\ {\mu}m$) are depicted for various eigenfrequencies ($\omega=-4.25$, $1.53$, $28.22$, and $141.15$). The color bar denotes the phase of the eigenvectors ($\theta={\tan}^{-1}(e_{y}/e_{x})$), capturing the directional information of the vibrational modes. The size of the arrow head indicates the magnitude of the eigenvector field at that point. Topological defects are indicated by filled circles, with red and blue representing $q=+1$ and $-1$ respectively, highlighting the presence of non-trivial topological features within the two-dimensional colloidal glass system.
@@
Figure 4: a The vibrational density of states ($D(\omega)$) and the total number of defects ($N_{d}$) are plotted against vibrational frequencies. The left-side vertical axis corresponds to $D(\omega)$, while the right-side axis corresponds to $N_{d}$. The inset illustrates the low-frequency behavior, with solid lines indicating linear trends. b Pearson correlation coefficients ($C$) between $D(\omega)$ and $N_{d}(\omega)$ are depicted for varying $\omega$, considering three different frequency widths: $\Delta\omega=50,75,$ and $100$. The shaded gray area marks the frequency range $\omega\in[250,340]$, where strong anti-correlation is observed in both panels a and b.

### v2

@@ Abstract
Topological defects are singularities within a field that cannot be removed by continuous transformations. The definition of these irregularities requires an ordered reference configuration, calling into question whether they exist in disordered materials, such as glasses. However, recent work suggests that well-defined topological defects emerge in the dynamics of glasses, even if they are not evident in the static configuration. In this study, we reveal the presence of topological defects in the vibrational eigenspace of a two-dimensional experimental colloidal glass. These defects strongly correlate with the vibrational features and spatially correlate with each other and structural “soft spots”, more prone to plastic flow. This work experimentally confirms the existence of topological defects in disordered systems revealing the complex interplay between topology, disorder, and dynamics.
@@ Introduction
Topological defects (TD) represent a ubiquitous hallmark of nature across different scales and they are generally defined as singularities in a local order parameter. They can appear in a wide range of physical systems [1, 2, 3], including but not limited to liquid crystals, superconductors, superfluids, ferromagnets, biological systems [4], and also our early universe [5]. Despite being microscopic in nature, TD can exert macroscopic influences on the behavior of the entire system. For example, they have a strong influence on the optical properties of liquid crystals and play crucial roles in various biological processes, including cell division, tissue formation, and the motion of cellular aggregates [6].
@@
In the realm of condensed matter physics, the emergence of these defects disrupts ordered states [7] influencing collective excitations and even triggering phase transitions, such as the Kosterlitz, Thouless, Halperin, Nelson, Young (KTHNY) transition [8] in two-dimensional solids, investigated e.g. in a colloidal monolayer [9]. These defects possess distinct (topological) charges and exhibit robustness against continuous structural deformations, rendering them essential for understanding the fundamental properties of materials and being totally different from other non-topological defects such as vacancies or interstitials.
@@
Topological defects in 3D crystalline solids manifest as dislocations and disclination lines [7], providing structural irregularities within the otherwise ordered lattice, and the elementary carriers of plasticity. Dislocations, in particular, allow for the motion of crystalline planes sliding over each other (glide motion). The plasticity of crystalline solids is by now fully understood in terms of dislocation dynamics and dislocation networks, based on seminal concepts introduced by Taylor [10], Polanyi [11] and Orowan [12]. In 2D, topological defects are point defects and their thermal dissociation causes the elastic moduli to disappear [13].
@@
In amorphous solids, instead, the absence of long-range order complicates the identification and characterization of topological defects (since, by definition, an ordered background is needed in order to detect its irregularities), posing a big challenge in linking structure with dynamics and predicting mechanical properties from the undeformed material configuration. The chase for structural topological defects in amorphous solids, borrowing from the concepts widely used for crystals, is certainly not new and dates back to the early 70’s [14]. Despite various efforts [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], it was nevertheless concluded that topological quantities cannot be properly defined by looking at the structural characteristics of glasses.
@@
Therefore, apart from isolated cases [27, 28], the identification of the plastic carriers and of the “soft spots”, the regions (or particle clusters) more prone to plastic flow, in glasses remained based until now on the so-called structural indicators [29] and on the phenomenological concept of shear transformation zones (STZ) [30, 31]. On the other hand, structural “soft spots”, where particle rearrangements are initiated, have been postulated to be analogous to dislocations in crystalline solids [32] but lacking any formal definition and topological character.
@@
Recent advances in simulation and theory [33] have pursued the idea of looking for topological defects in glasses using dynamical quantities such as the displacement vector field or the eigenvector field of normal mode vibrations, instead of the seemingly featureless structure. Following this idea, Baggioli et al. [34] have unveiled well-defined topological defects in polymer glasses under shear deformation. These topological defects are predicted by theory [35] as singularities in the microscopic displacement field, and are mathematically described by a continuous-valued Burgers vector (in accord with an early intuition of Kleman and Friedel [36]. In amorphous solids, the atomic displacement field under an external deformation comprises two vector fields, one of which is fully ordered and is called the affine displacement field. This is simply the trajectory along which each atom moves as a consequence of the external strain field. The other vector field is known as the nonaffine displacement field and is much more random and irregular. It arises due to the local force imbalance on each atom in its affine position, where it is subject to a non-vanishing force (due to the lack of inversion symmetry) communicated by its neighbours [37]. In this sense, the affine component of the displacement field represents the ordered background disrupted by the topological defects living within (but not being equivalent to) the nonaffine displacements.
@@
Even more recently, mixing the ideas of [34] and [38], numerical investigations by Falk and collaborators [39] have identified saddle-like topological defects with quantized -$1$ charge in the displacement field, generating Eshelby-like quadrupolar fields which align to form the shear bands responsible for the yielding of amorphous solids. Interestingly, shear banding is directly related to the percolation of STZ, whose dynamics have been shown [40] to be intimately connected to vortex-like structures as those proposed by Wu et al.[38]. Finally, although topological charge neutrality is always enforced by the conservation of global elastic dipole charge, the yielding process in amorphous systems has been related to the clustering of net negative topological charge [41].
Topological defects in crystals can be also defined in terms of geometrical properties (curvature and torsion) [7]. A similar approach based on geometry has been recently proposed by Moshe, Procaccia and collaborators as the origin of plastic screening in amorphous materials [42], where the defects are identified with the elastic (dipolar and quadrupolar) charges. Interesting results in this context can be also obtained in terms of a tensorial form of electromagnetism known as vector charge tensor gauge theory that is able to reproduce the structure of static stress correlations in granular materials [43].
@@
Colloidal glasses have been proved to be an excellent experimental setup to test the validity of various measures and theories related to structure and dynamics [44]. Using optical microscopy, one has direct access to the structural information at the particle level. In this experimental study, we use bidisperse super-paramagnetic colloidal particles which are confined by gravity at an atomically flat interface to form a two dimensional disordered structure. Applying an outer magnetic field perpendicular to the monolayer, the particles interact via well defined dipole-dipole repulsion. This gives a specific type of colloidal particles which become magnetised in the presence of a magnetic field and form a two-dimensional disordered structure. The particles interact via dipole-dipole repulsive interaction, so, based on particle positions, we can also directly infer the pairwise interaction energy. Using this information, we construct the dynamical matrix (Hessian) of the system, and perform the identification of topological defects in the field of normal modes. In this process, the analysis of the experimental data reveals unique features of vibrational characteristics of the system, and also the corresponding correlation with topology and “soft spots”. Our experimental study demonstrates the existence of well-defined topological defects in the vibrational field of a finite temperature 2D colloidal glass, suggesting further experimental studies that can employ these defects to understand the thermal and mechanical properties of many complex, disordered systems.
@@
Figure 1: Experimental realization of the colloidal monolayer forming a glass. a, Schematic showcasing the binary colloidal mixture with large and small particles colored in silver and pink (side view, not to scale). A water droplet hanging by surface tension inside a cylindrical hole in a glass plate, hosts the mixture of colloidal particles at the bottom water-air interface. The volume of water droplet is actively regulated via a computer controlled microsyringe to ensure a flat interface. The zoomed portion shows the dipole moment induced in each colloidal particle in the presence of external magnetic field $\vec{H}$ perpendicular to the monolayer that leads to repulsive dipolar interaction. b, Top view of the arrangement of colloidal particles, reconstructed from positional data by video microscopy.
Our experimental setup comprises a colloidal monolayer, where individual particles sediment to a flat water/air interface in a hanging-droplet configuration under the influence of gravity. These colloidal particles have two different masses and diameters, forming a binary mixture (Fig. 1b), and exhibit Brownian motion within the two-dimensional plane (see Supplementary Materials). The particles consist of polystyrene with incorporated nanograins of iron oxide, thus interactions are paramagnetic in nature, and a magnetic dipole moment is induced in each particle when an external magnetic field is applied perpendicular to the surface. The particles are sterically stabilized with Sodium Dodecyl Sulfate (SDS) slightly below critical micelle concentration. As anionic tenside, residual surface charges of the colloids are effectively shielded by counter-ions. The hard-core diameter (including the SDS layer) is never probed by $k_{B}T$ and the interaction is given by a pairwise dipole-dipole potential of the form [45, 46] $E_{\text{pot}}\sim 1/r^{3}$, where $r$ is the inter-particle distance. The strength of the interaction can be controlled by the external magnetic field $H$, setting the characteristic energy scale of the system. A crucial parameter governing the system’s behavior is represented by $\Gamma=E_{\text{pot}}/E_{\text{kin}}$, denoting the ratio of potential energy (stemming from this mutual dipolar interaction among particles) to the kinetic energy ($E_{\text{kin}}\sim T$) arising from thermal motion [47]. It can be interpreted as an inverse temperature or more didactically as dimensionless pressure. For our study, we set $\Gamma=423$, a value at which the system is deep in a glassy state [47]. Optical microscopy is employed to record the particle positions in the field of view at a certain interval of time (Fig. 1a). The experimental setup is described in detail in [48], while the structure and dynamics of this system have been studied in previous work [46, 45].
@@
Using the long-range dipole-dipole interaction potential among the particles, we numerically construct the dynamical (Hessian) matrix which is diagonalized to obtain the eigenvectors $\vec{e}_{l}$ and eigenfrequencies $\lambda_{l}$. Here, $\lambda_{l}$ indicates the l-th eigenvalue with $l=1,...,2N$; $N$ being the number of particles in the field of view. From the eigenvalues, we derive the corresponding eigenfrequencies $\omega_{l}^{2}=\lambda_{l}$ (using a mass-rescaled Hessian, see Methods). As expected from amorphous solids at finite temperature [49], the spectrum displays a fraction of negative eigenvalues, corresponding to unstable modes with purely imaginary frequency. We notice that the presence of unstable modes is not incompatible with the solid nature of our experimental system. In fact, unstable modes exist in glasses [49] and even heated crystals [50], as a result of local anharmonicities. More in general, these unstable modes correspond to dynamics over regions of the potential landscape with locally negative curvature (e.g., saddle points and potential barriers) and they are widely observed in supercooled liquids as well [51]. For the unstable part of the spectrum ($\lambda<0$), we follow the practice of re-defining a positive definite frequency $\tilde{\omega}=-i\sqrt{\lambda}$ and plotting the corresponding vDOS on the negative frequency axes upon identifying $\omega=-\tilde{\omega}$. The results for the experimental vibrational density of states (vDOS) $D(\omega)$ are shown in Fig. 2a.
@@
As highlighted in the inset of Fig. 2a, the low frequency behavior of the vDOS shows a linear scaling, $D(\omega)\sim\omega$. This behavior is compatible with Debye’s law in two spatial dimensions but it is also a characteristic feature of the vDOS of systems with unstable modes [52, 53]. Interestingly, we observe that the vDOS is symmetric at low frequency and the linear scaling (including the corresponding slope) is the same for the stable and unstable branches. We also notice that only the stable branch can be directly interpreted as pertaining to vibrational modes stricto sensu. We have further tested the robustness of these findings by introducing a cut-off distance $R_{c}$ in the interaction potential, incorporating random polydispersity in particle masses, and accounting for experimental measurement errors in distances (see Supplementary Figs. S3 and S4). Our results demonstrate that the characteristics of the vDOS remain robust against variations in $R_{c}$, polydispersity, and possible experimental measurement errors.
@@
In Fig. 2b, we present the same vDOS normalized by the low-frequency linear (Debye) scaling, $D(\omega)/\omega$. In this representation, the presence of an anomalous peak is clear around a characteristic frequency $\omega_{\rm BP}$, marked within error bars by the shaded vertical strip ($15-17$ Hz). This vibrational anomaly over Debye’s law is known as “boson peak” and it is a common feature of amorphous materials [54]. Despite the huge effort devoted to characterize and comprehend this feature in 3D systems, its experimental detection in quasi two-dimensional (2D) amorphous materials remains limited [55, 56]. In this respect, our result provides another experimental identification of the BP anomaly in a specific type of 2D amorphous materials – colloidal glasses – using directly the Hessian matrix rather than the covariance matrix [57]. In these colloidal systems the density of states using covariance matrix have been extensively studied previously [57, 47]. In anharmonic systems like ours, the vDOS obtained from the covariance matrix formalism differs significantly from that obtained through normal mode analysis (see, e.g., [58]). Therefore, a direct comparison between these two methods appears not trivial and requires further systematic investigations. Also, the unit of $\omega$ (derived from the normal mode analysis) is ${\textrm{sec}}^{-1}$, while the frequency unit obtained from the covariance matrix is $\mu m^{-1}$ [57, 47], making this comparison more difficult.
@@
To measure the extent of mode localization, we also calculate the participation ratio $P(\omega)$, plotted in Fig. 2c. We observe a sharp peak corresponding to acoustic phonon-type excitations at very low (positive) frequency, and a secondary broader peak at a higher frequency possibly linked to optical-like excitations due to the particle size mismatch. This secondary peak then drops sharply in correspondence of the Anderson-localized highest frequency modes which terminate at the Debye frequency of the system.
@@
Figure 2: Vibrational characteristics of the system. a, The vibrational density of states $D(\omega)$ is presented, with the inset highlighting the low-frequency behavior. The dashed line represents a linear trend, consistent with Debye’s prediction $D(\omega)\sim\omega^{d-1}$, where $d=2$. b, $D(\omega)/\omega$ versus $\omega$ is plotted at low frequencies, with shaded regions indicating the occurrence of the boson peak. c, The participation ratio $P(\omega)$ versus $\omega$ is depicted, providing insights into the localization properties of vibrational modes across different frequency ranges.
@@
Figs. 3a-d provide visual representations of the eigenvector fields at various frequencies, with filled red and blue circles denoting the locations of vortices and anti-vortices, respectively. Additionally, the color map indicates the local angle of the eigenvector field with respect to the $x$-axes. The structure of the eigenvector field is evidently highly heterogeneous, displaying several vortex-like structures and singularities. Notably, the eigenvectors at lower frequencies exhibit a smaller number of topological defects. This suggests that the number of defects grows with frequency and becomes more uniformly distributed in space. Significantly, at lower frequencies, there are fewer topological defects, and a rich, cooperative and swirling eigenfield structure. This might be related to the fact that at low-frequency the dynamics are dominated by plane waves, displaying a periodic structure of swirling (see e.g. Fig.3 in [59]). Conversely, at higher frequencies, there is a notable increase in defect density, with defects uniformly distributed throughout the space and reduced coherence in the eigenvector field. Interestingly, topological defects can be also found in the eigenvectors corresponding to unstable modes, e.g. panel a in Fig. 3.
@@
Figure 3: Topological defects in the eigenvector field of the glass. a-d, Eigenvector fields within a rectangular field of view ($54a\times 40a$, where $a=20.78\leavevmode\nobreak\ {\mu}m$) are depicted for various eigenfrequencies ($\omega=-4.25$, $1.53$, $28.22$, and $141.15$). The color bar denotes the phase of the eigenvectors ($\theta={\tan}^{-1}(e_{y}/e_{x})$), capturing the directional information of the vibrational modes. The size of the arrow head indicates the magnitude of the eigenvector field at that point. Topological defects are indicated by filled circles, with red and blue representing $q=+1$ and $-1$ respectively, highlighting the presence of non-trivial topological features within the two-dimensional colloidal glass system.
Fig. 4a presents the vibrational density of states $D(\omega)$ and the total number of defects $N_{d}$ against vibrational frequencies within the same frame. This highlights a robust correlation between these quantities across the entire frequency spectrum. The inset of Fig. 4a delves into the low-frequency behavior of $N_{d}$, where the dashed line signifies the linear dependence on $\omega$. To further establish the correlation between vibrational properties and topological defects, we calculate the Pearson correlation coefficient, $C(\omega)$, at various $\omega$ values, considering data points within a specified frequency range $\omega-\Delta\omega/2$ to $\omega+\Delta\omega/2$ (see Supplementary Materials). Fig. 4b illustrates the variation of $C(\omega)$ for different frequency width values ($\Delta\omega=50$, $75$, and $100$). We observe that the fluctuations of $C(\omega)$ decrease as the frequency width increases. The proximity of $C$ values to unity across the majority of the spectrum signifies a robust correlation between $D(\omega)$ and $N_{d}$. However, intriguingly, instances of strong anti-correlation are observed within specific frequency ranges, as highlighted by gray vertical stripes in Fig. 4a and Fig. 4b. Particularly noteworthy is the presence of such anti-correlation within the frequency range $\omega\in[250,340]$, prompting further theoretical investigation to unravel its origin and its consequential impact on the system’s structure and dynamics. This observed anti-correlation arises because, in this frequency range, the DOS is decreasing while the number of defects continues to increase gradually. At this moment, we do not have a concrete understanding of the physical significance of these anti-correlations appearing at large frequencies, above the Debye regime where both $N_{d}(\omega)$ and $D(\omega)$ scale like $\omega$ (and nicely correlate).
It is important to notice that our scaling $D(\omega)\propto\omega$ at low frequency seems at first sight in contradiction with the results reported in Ref. [38]. In order to reveal the cause of this discrepancy, we emphasize that the results of ref. [38] are obtained for a zero temperature system. Here, we perform additional simulations in a 2D Lennard-Jones glass at zero as well as finite temperatures (see Supplementary Information for details). At $T=0$, we reproduce the results in Ref. [38], where $N_{d}$ scales quadratically with the frequency. We further extend our analysis at finite temperatures. As expected, the fraction of unstable modes grows with temperature, in contrast to the athermal case where such modes are absent, as shown in 5a. Our focus is on the small $\omega$ behavior of $N_{d}$ with the temperature change. In Fig. 5b we observe that, unlike for the athermal case where $N_{d}\sim\omega^{\alpha}$ with $\alpha=2$, the scaling of the number of defects with frequency becomes linear at finite temperature. The scaling of the number of defects in the experimental system shows an exponent $\alpha$ close to $1$, compatible with the simulation results at finite $T$. In Fig. 5b, we show that this difference is due to the finite temperature dissipative effects and we recover the results of [38] in the zero temperature limit, solving the apparent contradiction.
@@
Figure 4: Correlation between vibrational density of states and number of topological defects. a, The vibrational density of states ($D(\omega)$) and the total number of defects ($N_{d}$) are plotted against vibrational frequencies. The left-side vertical axis corresponds to $D(\omega)$, while the right-side axis corresponds to $N_{d}$. The inset illustrates the low-frequency behavior, with dashed line indicating linear trend. b, Pearson correlation coefficients ($C$) between $D(\omega)$ and $N_{d}(\omega)$ are depicted for varying $\omega$, considering three different frequency widths: $\Delta\omega=50,75,$ and $100$. The shaded gray area marks the frequency range $\omega\in[250,340]$, where strong anti-correlation is observed in both panels a and b.
Figure 5: Simulation of 2D Lennard-Jonnes glass [38] at finite temperature. a, Vibrational density of states $D(\omega)$ and b, number of topological defects $N_{d}$ as a function of frequency $\omega$ are shown in linear and double-log scale, respectively, at different temperatures for a two dimensional Lennard-Jones glass.